solve by completing the square:
x^2-22x-4=0

Answer by mananth(16946)

x^2-22x-4
x^2-22x+121-4 =121
x^2-22x+121=121+4
(x-11)^2=125
(x-11)^2=25*(5)
Take the square root on both sides
x-11= +/- 5sqrt(5)
x=11+5sqrt(5)
OR
x=11-5sqrt(5)